OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example 7(sqrt)a can be written as 7(sqrt)a.

OpenStudy (anonymous):

Hey welcome to OpenStudy. Make sure you have read the code of conduct. http://openstudy.com/code-of-conduct Now to your question, Hope that this will help you understanding - so radical indice 2 sign squareroot what is equal in exponentially form by 1/2 - radical indice 3 is cuberoot what in exponent form is 1/3 - so for your help for example 4^(1/2 = (2^2)^(1/2) =2^(2*1/2) = =2^(2/2) =2^1 =2 - but this is squareroot of 4 =2 - ok ?